OpenStudy (anonymous):

find two mpositive real numbers whose product is a maximum and whose sum of the first number and four times the second is 240

OpenStudy (dumbcow):

A + 4B = 240 A = 240 - 4B A*B = (240-4B)*B = -4B^2 + 240B Use formula for finding line of symmetry of this parabola: B = -240/(2*-4) = -240/-8 = 30 This value for B maximizes A*B. By substitution A = 240-(4*30) = 240 - 120 = 120 Therefore A = 120, B = 30